OpenStudy (anonymous):

Can someone explain how to find the derivative of -16t^2+13t+1?

OpenStudy (abb0t):

Use the fact that: \[\frac{ d }{ dx }(x^n) = n x^{n-1}\] and simialrly: \[\frac{ d }{ dx }(cx^n)= ncx^{n-1}\]